Three individuals form a partnership and agree to divide the profits equally. X invests $9,000, Y invests $7,000, and z invests $ 4,000. If the profits are $4.800, how much less does X receive than if the profits were divided in proportion to the amount invested?

Answer by htmentor(1343)

If divided equally, the profits are 4800/3 = 1600 for each investor.
X's portion of the investment is 9000/(9000+7000+4000) = 9/20
So if divided in proportion to his investment, X would have received 9/20*4800 = 2160.
So he receives 2160 - 1600 = $560 less.